OpenStudy (anonymous):

Please check my answer if I am wrong I will redo it. solve: sqrtx+14-12=x x = –10 x = –13 x = –10 and x = –13 None of the above I got C am I correct

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (anonymous):

UNICORNS ARE NOT REAL

terenzreignz (terenzreignz):

Ambiguity is fatal.

OpenStudy (anonymous):

i kniw unicorns are not real i just like them

OpenStudy (anonymous):

is my answer correct or do i need to redo it

terenzreignz (terenzreignz):

Is it this: \[\Large \sqrt{x +14 }-12 = x\]

OpenStudy (anonymous):

Simplify both sides, then kill the uniciorn

OpenStudy (anonymous):

yes that is my problem and stop picking on my unicorn @Jack17

OpenStudy (anonymous):

send the 12 other side and square both sides

OpenStudy (anonymous):

\(x+14 = (x+12)^2 \)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Its a quadratic equation with integer coeiffients

OpenStudy (anonymous):

and a dead unicorn

OpenStudy (anonymous):

haha very funny unicorns even though are not real i love them

OpenStudy (anonymous):

just by looking at the options, and equation x=-10 satisfies the equation x= -13 doesnt satisfy the equation so go with option A

OpenStudy (anonymous):

ok thanks that is what i got when i redid it

OpenStudy (anonymous):

good :)